i have a couple questions
what is the following answers?
(17)5x-7=2x+8
(18)2y+3=y+5

Answer by checkley71(8403)

5x-7=2x+8
5x-2x=8+7
3x=15
x=15/3
x=5
proof
5*5-7=2*5+8
25-7=10+8
18=18
------------------------------------------
2y+3=y+5
2y-y=5-3
y=2
proof
2*2+3=2+5
4+3=7
7=7